Alice borrows $16,000 to buy a car. She pays 5.4% interest compounded monthly and the loan is for 2 years. Find the monthly payments. (Round your answer to two decimal places.)
$

Answer by Boreal(15235)

Principal*(r/n)=Payment (1-(1+(r/n))^-nt
16000*(.054/12)=payment(1-1+(.054/12))^24
72=P*(0.10215) and do without rounding until the end.
$704.81